OpenStudy (anonymous):

HELP PLEASE!!!!!!!!! The volume of a cube is given by V=s^3, where s is the length of a side. Find the length of a side of a cube if the volume is 800 cm^3

jimthompson5910 (jim_thompson5910):

V = s^3 ... given formula 800 = s^3 ... plug in the given volume V = 800 s^3 = 800 s = ???

OpenStudy (anonymous):

@hartnn can you help me. I am still not understanding this

jimthompson5910 (jim_thompson5910):

the next step would be to take the cube root of both sides

OpenStudy (anonymous):

there is nothing cubed that gives you 800

jimthompson5910 (jim_thompson5910):

so it's going to be some decimal number

OpenStudy (anonymous):

that is what I have been trying to figure out lol

jimthompson5910 (jim_thompson5910):

that's why you need a calculator

OpenStudy (anonymous):

thats what i am doing

jimthompson5910 (jim_thompson5910):

cube root of 800 = 800^(1/3) = ???

OpenStudy (anonymous):

i got 9.283^3

jimthompson5910 (jim_thompson5910):

me too, so the side is roughly 9.283 cm

OpenStudy (anonymous):

the part the got me was that the volume = 800cm^3

OpenStudy (anonymous):

i mean that

jimthompson5910 (jim_thompson5910):

if it was 8 or 8000, then it would have been a whole number since 2^3 = 8 and 20^3 = 8000 but it's perfectly possible to get some length that's not a whole number

OpenStudy (anonymous):

ok thank you for your help

jimthompson5910 (jim_thompson5910):

np